What is the genral formula for converting the X,Y co-ordinates in Radius,Theta system?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

these are different in different cases ...

 

in case of ellipse (X,Y) = (acos@,bsin@)                     X2/a2 + Y2/b2 = 1

 

in case of circle (X,Y) = (Rcos@,Rsin@)                (X2+Y2 = R2)

 

in case of hyperbola (X,Y) = (asec@,btan@)                X2/a2 - Y2/b2 = 1

 

these are chosen in such a way so that it can satisfy the curve ....